# Both Stellar Mass and Gravitational Potential Shape the Gas-Phase Metallicity

### v1

@@ 3.1 Partial Correlation Coefficients
To investigate the direct relationships between variables while accounting for the influence of others, we perform a Partial Correlation Coefficient (PCC) analysis, following a methodology similar to that described in Baker et al. (2022b) and Baker & Maiolino (2023). This approach allows us to quantify the strength of the association between two variables while statistically removing the effect of other variables. By doing so, we can isolate genuine correlations and distinguish them from those that arise indirectly through shared dependencies.
@@
In our analysis, we conceptualise the three variables within a three-dimensional framework: $\mathrm{SFR}$ (y-axis) plotted against $\mathrm{M_{*}}$ (x-axis), with gas-phase metallicity visualised through a third dimension (e.g., colour-coding, see also Bluck et al., 2020a; Piotrowska et al., 2020; Baker et al., 2022a, b; Baker & Maiolino, 2023; Baker et al., 2024). Here, metallicity corresponds to variable A, $\mathrm{SFR}$ to B, and $\mathrm{M_{*}}$ to C.
To interpret the PCC results, we adopt the arrow angle formalism introduced by Bluck et al. (2020b), which identifies the direction of the steepest average increase in metallicity. The angle $\Theta$, measured from the horizontal axis (three o’clock position), is given by (Bluck et al., 2020b):
@@
To obtain arrow angle uncertainties, we applied the same method as Baker et al. (2022b) and Baker & Maiolino (2023) and utilised 100 bootstrap random samples and computed their standard deviation.
@@ 3.2 Random Forest
While PCCs are good for determining the direction of correlations, they do not apply to a wide variety of relations, as they only work for monotonic relations. Thus, we also employ a random forest Regression (Breiman, 2001) to more accurately determine the relative strengths of the correlations. This machine learning approach has previously been demonstrated to be effective in identifying fundamental relationships among galaxy properties (e.g. Baker et al., 2022b; Bluck et al., 2022; Baker & Maiolino, 2023; Sánchez-Menguiano et al., 2024a).
A random forest algorithm works as follows: it makes predictions by combining the results of many decision trees, each trained on different subsets of the data. Averaging the outputs of these trees improves accuracy and reduces the risk of overfitting to noise in the data. Thus, it is advantageous for our analysis as it allows us to simultaneously investigate several intercorrelated parameters. As was found by Bluck et al. (2020b, 2022); Baker & Maiolino (2023), the random forest Regressor can distinguish between intrinsic dependences and intercorrelated parameters. In our analysis, we utilise the random forest Regressor to determine the relative importance of different parameters driving the gas-phase metallicity.
We base our implementation of the algorithm on the scikit-learn Python package (Pedregosa et al., 2011). First, we split the data into a train and test set. For the MaNGA sample, we define our test set to be $30\%$ of the data, while for SDSS, we consider half of the selected sample. The training set consists of the predictor variables (all analysed galaxy parameters except the gas-phase metallicity) paired with the target variable (the gas-phase metallicity). The model is fit on this training data, and the test set is used to evaluate its predictive performance and derive the relative importance of the predictors in determining the target quantity. The galaxy properties of interest are the input predictors of the algorithm, and the gas-phase metallicity acts as our target feature. We rely on mean squared error (MSE) measurements of the training and test samples to determine the model’s precision. For example, suppose the MSE of the train sample is significantly less than the MSE of the test sample. In that case, we could be overfitting our data, i.e., including noise in finding the parameter importance. On the contrary, if the train MSE is much larger than the test MSE, we could be underfitting our data, meaning that the model is not exploring the full data space when computing the relative importance. To avoid overfitting and underfitting, we fine-tune the hyperparameters of our random forest regressor and use the following hyperparameters universally for both samples: $\mathrm{n_{estimators}=300,max_{depth}=30,min_{split}=2,}$and $\mathrm{min_{samples\_leaf}=6}$.
@@ 4.1 Dominant Features via the Random Forest
4.1.1 Metallicities within $1R_{e}$: the dominance of M∗,phot
@@
We start by simultaneously exploring, from our random forest analysis, the relative importance of multiple galactic properties in predicting the average gas-phase metallicity within $\mathrm{1R_{e}}$. First, we consider our selected MaNGA sample and analyse how different parameters (see table 2 for a description) influence the gas-phase metallicity. We include different stellar mass estimates. The spectroscopic method (denoted as $\mathrm{M_{*,spec}}$) consists of the sum of the spectroscopically determined masses in all spaxels of the MaNGA maps (or all spaxels within $\rm 1R_{e}$, in which case it is denoted as $\rm M_{*,spec,R_{e}}$); these are either measured ourselves by taking the sum of spaxels within a certain radius, or taken from the pyPipe3D catalogue. We also use the stellar masses provided in the MPA-JHU and pyPipe3D catalogue, which are determined from the central (single-fibre) SDSS spectroscopic information and corrected with photometry for the extended emission outside the fibre; this is denoted as $\mathrm{M_{*,phot}}$ (although it also has a spectroscopic component). Additionally, similar to previous studies (D’Eugenio et al., 2018; Barone et al., 2020; Vaughan et al., 2022; Cappellari, 2023; Baker & Maiolino, 2023; Sánchez-Menguiano et al., 2024a), we also include the “stellar gravitational potential” inferred via the stellar mass $\mathrm{\Phi_{*}=M_{*}/R_{e}}$, where $\mathrm{M_{*}}$ is replaced by the different mass estimates, in our analysis. This can be considered a proxy for the total gravitational potential in cases (and in apertures) where stars dominate the gravitational potential. We also include in the random forest analysis the $\mathrm{SFR}$ and effective radius $\mathrm{R_{e}}$. The former needs to be taken into account due to the metallicity’s secondary dependence, defined as the Fundamental Metallicity Relation (FMR, Mannucci et al., 2010). From Zhu et al. (2023) we include several dynamically inferred properties, namely the effective velocity dispersion $\mathrm{\sigma_{e}}$ and dynamical mass within the effective radius $\mathrm{M_{dyn,Re}}$, with which we also compute the dynamically inferred gravitational potential within $\rm 1R_{e}$, $\mathrm{\Phi_{dyn,Re}=M_{dyn,Re}/R_{e}}$. Lastly, we include a uniform random variable as a control measure.
@@
The results on the relative importance of these parameters in determining the metallicity within $\rm 1R_{e}$ are shown in Figure 1, in the form of a bar chart. The stellar mass estimated via spectro-photometric measurements is the main driver of the metallicity. Stellar mass being the main driver of the gas-phase metallicity was also confirmed by Baker & Maiolino (2023), although for the spectroscopic measurement of the stellar mass (their analysis did not include the photometric stellar mass estimates) and for a dataset that is only about a quarter the size of ours. There is a secondary dependence of $\mathrm{\sim 17\%}$ on the stellar gravitational potential inferred via the spectro-photometric stellar mass, but a significantly larger error bar also accompanies it. Stellar mass and gravitational potentials inferred from pure spectroscopy are ruled out, and the dynamical properties (within 1Re) also have insignificant relative importance, eliminating them as possible drivers for the gas-phase metallicity within $\rm 1R_{e}$.
@@
Spectro-photometric estimates of stellar mass and gravitational potential overpowering estimates from IFU spectroscopy were also confirmed by Sánchez-Menguiano et al. (2024a). This likely reflects the fact that photometric masses, derived from broadband SED fitting, typically achieve higher S/N than spectroscopic masses, which rely on narrower spectral features that can be more sensitive to noise and template mismatches (Kauffmann et al., 2003; Drory et al., 2004). In addition, while photometric masses may be less accurate in an absolute sense due to modelling assumptions and systematic biases, they tend to be more precise, yielding consistent results across the galaxy population (Sánchez-Menguiano et al., 2024a). Tree-based algorithms like random forests naturally favour features with lower noise and higher variance, which may explain why photometric masses are given greater importance in our analysis. However, in contrast to our results, Sánchez-Menguiano et al. (2024a) find that $\mathrm{\Phi_{,phot}}$ ranks significantly higher than $\mathrm{M_{,phot}}$.
@@
Other previous studies also found a strong dependence on $\mathrm{\Phi_{*}}$ (D’Eugenio et al., 2018; Cappellari, 2023; Sánchez-Menguiano et al., 2024a). Generally, this importance is explained by $\mathrm{\Phi_{*}}$ being a tracer of the gravitational potential of the galaxy and being connected to retaining metals. However, we also include $\mathrm{\Phi_{dyn,Re}=M_{dyn}/R_{e}}$ in our analysis, which should give a more accurate measure of the gravitational potential than $\mathrm{\Phi_{*}}$. Yet, $\mathrm{\Phi_{dyn,Re}}$ has a negligible low relative importance according to our random forest analysis shown in Figure 1. We further investigate the correlation between $\mathrm{M_{dyn,R_{e}}}$ and the gas-phase metallicity via a differential uncertainty measurement test in Appendix D.
@@ 4.1.2 Radial dependence: $\rm M_{*}$ overthrown by $\Phi_{*}$ at large radii
Prompted by the puzzling discrepancies between different studies on the relative importance of stellar mass and gravitational potential in determining the gas metallicity, we have explored the role of the spatial scale on which such importance is assessed. More specifically, we extended the analysis performed in the previous section, which was on the metallicities estimated within $\rm 1R_{e}$, by re-running our random forest regression for metallicities computed within different radii, $\in[0.5,1.0,1.5,2.0]$ in units of $\rm R_{e}$.
@@
Since we have already established that the stellar mass and gravitational potential from photometry, and the SFR are the three most important drivers of the gas-phase metallicity, we solely focus on these parameters. The results are shown in Figure 2. It is evident that both $\mathrm{M_{*}}$ and $\Phi_{*}$ have a radial dependence. At small radii, the stellar mass is by far the dominant quantity determining the metallicity. As the radius increases, the metallicity is less impacted by changes in the stellar mass but increasingly correlates with the gravitational potential. We also include SFRs computed at different radii and the global SFR taken from the pyPipe3D catalogue in our analysis, but find no significant difference in the role of SFR as a function of radius.
@@
The most likely interpretation is that in the inner regions of galaxies, the strong correlation between metallicity and stellar mass arises because stellar mass traces the integrated metal production from the star formation history (Sánchez et al., 2014; Peng et al., 2015; Chisholm et al., 2018; Maiolino & Mannucci, 2019). Central regions, characterised by deep gravitational potentials and high stellar densities, have short chemical enrichment timescales and efficient recycling of metals through stellar feedback, enabling evolution with little metal loss regardless of the overall gravitational potential (Baker & Maiolino, 2023).
@@
In the outer regions, where metals are less gravitationally bound, the gravitational potential (of which $\Phi_{*}$ is a proxy) becomes more important in regulating gas-phase metallicity. These regions are more susceptible to gas inflows, feedback-driven outflows, and radial migration, all of which can dilute or redistribute metals (Fraternali, 2017; Christensen et al., 2016; Tumlinson et al., 2017). The longer chemical mixing and cycling timescales in the outskirts (Ho et al., 2015; Tissera et al., 2019) mean they may not reach chemical equilibrium, making gravitational potential a better predictor of local metallicity than stellar mass.
@@
We finally note that the importance of dynamical gravitational potential ($\rm\Phi_{dyn,R_{e}}$) is minimal, and remains so regardless of the radius within which the metallicity is estimated. This is likely because the gravitational potential was estimated by Zhu et al. (2023) within $\rm 1R_{e}$, in the region where stellar mass dominates the importance. Unfortunately, dynamical masses estimated at larger radii are not available.
@@ 4.1.3 The relative importance in integrated SDSS spectra
Figure 2: Importances in driving the gas-phase metallicity at different radii, as indicated by the x-axis, of the photometric stellar mass $\mathrm{M_{*,phot}}$, photometric baryonic gravitational potential $\mathrm{\Phi_{*,phot}}$, global $\mathrm{SFR}$ from all star-forming spaxels from the pyPipe3D catalogue (Sánchez et al., 2022), and integrated $\mathrm{SFR}$ obtained by taking the sum of selected star-forming spaxels within different radii. The results indicate a spatial variation: at smaller radii, the stellar mass dominates over the baryonic gravitational potential, but this flips toward larger radii ($\mathrm{\gtrsim 1.25R_{e}}$).

### v2

@@ 3.1 Partial Correlation Coefficients
To investigate the direct relationships between variables while accounting for the influence of others, we perform a Partial Correlation Coefficient (PCC) analysis, following a methodology similar to that described in baker_metallicitys_2022 and baker_stellar_2023. This approach allows us to quantify the strength of the association between two variables while statistically removing the effect of other variables. By doing so, we can isolate genuine correlations and distinguish them from those that arise indirectly through shared dependencies.
@@
In our analysis, we conceptualise the three variables within a three-dimensional framework: $\mathrm{SFR}$ (y-axis) plotted against $\mathrm{M_{*}}$ (x-axis), with gas-phase metallicity visualised through a third dimension (e.g., colour-coding, see also bluck_are_2020; piotrowska_towards_2020; baker_almaquest_2022; baker_metallicitys_2022; baker_stellar_2023; baker_different_2024). Here, metallicity corresponds to variable A, $\mathrm{SFR}$ to B, and $\mathrm{M_{*}}$ to C.
To interpret the PCC results, we adopt the arrow angle formalism introduced by bluck_how_2020, which identifies the direction of the steepest average increase in metallicity. The angle $\Theta$, measured from the horizontal axis (three o’clock position), is given by (bluck_how_2020):
@@
To obtain arrow angle uncertainties, we applied the same method as baker_metallicitys_2022 and baker_stellar_2023 and utilised 100 bootstrap random samples and computed their standard deviation.
@@ 3.2 Random Forest
While PCCs are good for determining the direction of correlations, they do not apply to a wide variety of relations, as they only work for monotonic relations. Thus, we also employ a random forest Regression (breiman_random_2001) to more accurately determine the relative strengths of the correlations. This machine learning approach has previously been demonstrated to be effective in identifying fundamental relationships among galaxy properties (e.g. baker_metallicitys_2022; bluck_quenching_2022; baker_stellar_2023; sanchez-menguiano_stellar_2024).
A random forest algorithm works as follows: it makes predictions by combining the results of many decision trees, each trained on different subsets of the data. Averaging the outputs of these trees improves accuracy and reduces the risk of overfitting to noise in the data. Thus, it is advantageous for our analysis as it allows us to simultaneously investigate several intercorrelated parameters. As was found by bluck_how_2020; bluck_quenching_2022; baker_stellar_2023, the random forest Regressor can distinguish between intrinsic dependences and intercorrelated parameters. In our analysis, we utilise the random forest Regressor to determine the relative importance of different parameters driving the gas-phase metallicity.
We base our implementation of the algorithm on the scikit-learn Python package (pedregosa_scikit-learn_2011). First, we split the data into a train and test set. For the MaNGA sample, we define our test set to be $30\%$ of the data, while for SDSS, we consider half of the selected sample. The training set consists of the predictor variables (all analysed galaxy parameters except the gas-phase metallicity) paired with the target variable (the gas-phase metallicity). The model is fit on this training data, and the test set is used to evaluate its predictive performance and derive the relative importance of the predictors in determining the target quantity. The galaxy properties of interest are the input predictors of the algorithm, and the gas-phase metallicity acts as our target feature. We rely on mean squared error (MSE) measurements of the training and test samples to determine the model’s precision. For example, suppose the MSE of the train sample is significantly less than the MSE of the test sample. In that case, we could be overfitting our data, i.e., including noise in finding the parameter importance. On the contrary, if the train MSE is much larger than the test MSE, we could be underfitting our data, meaning that the model is not exploring the full data space when computing the relative importance. To avoid overfitting and underfitting, we fine-tune the hyperparameters of our random forest regressor and use the following hyperparameters universally for both samples: $\mathrm{n_{estimators}=300,max_{depth}=30,min_{split}=2,}$and $\mathrm{min_{samples\_leaf}=6}$.
@@ 4.1 Dominant Features via the Random Forest
4.1.1 Metallicities within $\rm 1R_{e}$: the dominance of $\rm M_{*,phot}$
@@
We start by simultaneously exploring, from our random forest analysis, the relative importance of multiple galactic properties in predicting the average gas-phase metallicity within $\mathrm{1R_{e}}$. First, we consider our selected MaNGA sample and analyse how different parameters (see table 2 for a description) influence the gas-phase metallicity. We include different stellar mass estimates. The spectroscopic method (denoted as $\mathrm{M_{*,spec}}$) consists of the sum of the spectroscopically determined masses in all spaxels of the MaNGA maps (or all spaxels within $\rm 1R_{e}$, in which case it is denoted as $\rm M_{*,spec,R_{e}}$); these are either measured ourselves by taking the sum of spaxels within a certain radius, or taken from the pyPipe3D catalogue. We also use the stellar masses provided in the MPA-JHU and pyPipe3D catalogue, which are determined from the central (single-fibre) SDSS spectroscopic information and corrected with photometry for the extended emission outside the fibre; this is denoted as $\mathrm{M_{*,phot}}$ (although it also has a spectroscopic component). Additionally, similar to previous studies (deugenio_gas-phase_2018; barone_gravitational_2020; vaughan_sami_2022; cappellari_full_2023; baker_stellar_2023; sanchez-menguiano_stellar_2024), we also include the “stellar gravitational potential” inferred via the stellar mass $\mathrm{\Phi_{*}=M_{*}/R_{e}}$, where $\mathrm{M_{*}}$ is replaced by the different mass estimates, in our analysis. This can be considered a proxy for the total gravitational potential in cases (and in apertures) where stars dominate the gravitational potential. We also include in the random forest analysis the $\mathrm{SFR}$ and effective radius $\mathrm{R_{e}}$. The former needs to be taken into account due to the metallicity’s secondary dependence, defined as the Fundamental Metallicity Relation (FMR, mannucci_fundamental_2010). From zhu_manga_2023 we include several dynamically inferred properties, namely the effective velocity dispersion $\mathrm{\sigma_{e}}$ and dynamical mass within the effective radius $\mathrm{M_{dyn,Re}}$, with which we also compute the dynamically inferred gravitational potential within $\rm 1R_{e}$, $\mathrm{\Phi_{dyn,Re}=M_{dyn,Re}/R_{e}}$. Lastly, we include a uniform random variable as a control measure.
Figure 2: Importances in driving the gas-phase metallicity in annuli (of width $\rm 0.2R_{e}$), as a function of Galactocentric radius of the annulus, of the photometric stellar mass $\mathrm{M_{*,phot}}$, photometric baryonic gravitational potential $\mathrm{\Phi_{*,phot}}$, global $\mathrm{SFR}$ from all star-forming spaxels from the pyPipe3D catalogue (sanchez_sdss-iv_2022), and integrated $\mathrm{SFR}$ obtained by taking the sum of selected star-forming spaxels within the same annuli. The x-axis corresponds to the radial position of the centre of each $\rm 0.2R_{e}$ annulus. The results indicate a spatial variation: at smaller radii, the stellar mass dominates over the baryonic gravitational potential, but this flips toward larger radii ($\mathrm{\gtrsim 0.9R_{e}}$).
@@
The results on the relative importance of these parameters in determining the metallicity within $\rm 1R_{e}$ are shown in Figure 1, in the form of a bar chart. The stellar mass estimated via spectro-photometric measurements is the main driver of the metallicity. Stellar mass being the main driver of the gas-phase metallicity was also confirmed by baker_stellar_2023, although for the spectroscopic measurement of the stellar mass (their analysis did not include the photometric stellar mass estimates) and for a dataset that is only about a quarter the size of ours. There is a secondary dependence of $\mathrm{\sim 17\%}$ on the stellar gravitational potential inferred via the spectro-photometric stellar mass, but a significantly larger error bar also accompanies it. Stellar mass and gravitational potentials inferred from pure spectroscopy are ruled out, and the dynamical properties (within 1Re) also have insignificant relative importance, eliminating them as possible drivers for the gas-phase metallicity within $\rm 1R_{e}$.
@@
Spectro-photometric estimates of stellar mass and gravitational potential overpowering estimates from IFU spectroscopy were also confirmed by sanchez-menguiano_stellar_2024. This likely reflects the fact that photometric masses, derived from broadband SED fitting, typically achieve higher S/N than spectroscopic masses, which rely on narrower spectral features that can be more sensitive to noise and template mismatches (kauffmann_host_2003; drory_comparing_2004). In addition, while photometric masses may be less accurate in an absolute sense due to modelling assumptions and systematic biases, they tend to be more precise, yielding consistent results across the galaxy population (sanchez-menguiano_stellar_2024). Tree-based algorithms like random forests naturally favour features with lower noise and higher variance, which may explain why photometric masses are given greater importance in our analysis. However, in contrast to our results, sanchez-menguiano_stellar_2024 find that $\mathrm{\Phi_{,phot}}$ ranks significantly higher than $\mathrm{M_{,phot}}$.
@@
Other previous studies also found a strong dependence on $\mathrm{\Phi_{*}}$ (deugenio_gas-phase_2018; cappellari_full_2023; sanchez-menguiano_stellar_2024). Generally, this importance is explained by $\mathrm{\Phi_{*}}$ being a tracer of the gravitational potential of the galaxy and being connected to retaining metals. However, we also include $\mathrm{\Phi_{dyn,Re}=M_{dyn}/R_{e}}$ in our analysis, which should give a more accurate measure of the gravitational potential than $\mathrm{\Phi_{*}}$. Yet, $\mathrm{\Phi_{dyn,Re}}$ has a negligible low relative importance according to our random forest analysis shown in Figure 1. We further investigate the correlation between $\mathrm{M_{dyn,R_{e}}}$ and the gas-phase metallicity via a differential uncertainty measurement test in Appendix C.
@@ 4.1.2 Radial dependence: $\rm M_{*}$ overthrown by $\Phi_{*}$ at large radii
Prompted by the puzzling discrepancies between different studies on the relative importance of stellar mass and gravitational potential in determining the gas metallicity, we have explored the role of the spatial scale on which such importance is assessed. More specifically, we extended the analysis performed in the previous section, which was on the metallicities estimated within $\rm 1R_{e}$, by re-running our random forest regression for average metallicities within narrow annuli of width $\rm 0.2R_{e}$ from the centre of each galaxy out to $\rm 2R_{e}$.
@@
Since we have already established that the stellar mass and gravitational potential from photometry, and the SFR are the three most important drivers of the gas-phase metallicity, we solely focus on these parameters. The results are shown in Figure 2. It is evident that both $\mathrm{M_{*}}$ and $\Phi_{*}$ have a radial dependence. At small radii, the stellar mass is by far the dominant quantity determining the metallicity. As the radius increases, the metallicity is less impacted by changes in the stellar mass but increasingly correlates with the gravitational potential. We also include SFRs computed within the annuli and the global SFR taken from the pyPipe3D catalogue in our analysis, but find no significant difference in the role of SFR as a function of radius.
@@
The most likely interpretation is that in the inner regions of galaxies, the strong correlation between metallicity and stellar mass arises because stellar mass traces the integrated metal production from the star formation history (sanchez_characteristic_2014; peng_strangulation_2015; chisholm_metal-enriched_2018; maiolino_re_2019). Central regions, characterised by deep gravitational potentials and high stellar densities, have short chemical enrichment timescales and efficient recycling of metals through stellar feedback, enabling evolution with little metal loss regardless of the overall gravitational potential (baker_stellar_2023).
@@
In the outer regions, where metals are less gravitationally bound, the gravitational potential (of which $\Phi_{*}$ is a proxy) becomes more important in regulating gas-phase metallicity. These regions are more susceptible to gas inflows, feedback-driven outflows, and radial migration, all of which can dilute or redistribute metals (fraternali_gas_2017; christensen_-n-out_2016; tumlinson_circumgalactic_2017). The longer chemical mixing and cycling timescales in the outskirts (ho_metallicity_2015; tissera_oxygen_2019) mean they may not reach chemical equilibrium, making gravitational potential a better predictor of local metallicity than stellar mass.
@@
We finally note that the importance of dynamical gravitational potential ($\rm\Phi_{dyn,R_{e}}$) is minimal, and remains so regardless of the radius within which the metallicity is estimated. This is likely because the gravitational potential was estimated by zhu_manga_2023 within $\rm 1R_{e}$, in the region where stellar mass dominates the importance. Unfortunately, dynamical masses estimated at larger radii are not available.